# Beyond Labeling Oracles: What does it mean to steal ML models?

### v1

@@ 4.1 Experimental Setting
We evaluate ME attacks on both vision and NLP tasks. We measure the attacker’s performance as the accuracy difference between the victim and the attacker on the original task. We define a successful attack as one that (i) produces a model that performs on a task nearly as well as the victim model and (ii) achieves this with a lower or equivalent sample complexity to that of training the victim model. Unless stated otherwise, in all of our experiments we assume that the attacker has a black-box access to the victim model: the adversary issues a query, and the victim model responds with a vector that indicates the probability of classifying the input in each of the task classes. We note that this setting provides the most information to the adversary, and we will explore more limited settings in which the adversary is limited to label-only access. Benchmarks are compatible with the ME literature.
@@
4.2 Benchmarking model extraction and training
First, we evaluate our baseline attacker. Here, we assume the attacker has access to a randomly sampled subset of the victim’s true training dataset, and uses this data only to query the victim and train the attack model. Figure 2 shows the attack accuracy as a function of the percentage of the data available to the attacker. Since ME quality is mainly affected by the quality of the attacker queries, real victim training data represents one of the best query sets for the attacker (compressed datasets e.g. (Wang et al., 2018) or disjoint IND data can potentially lead to an equally capable extraction set). Figure 2 shows that with only $10\%$ of the data, the attacker already gets relatively high performance, i.e. $79.21\%$ out of $95.54\%$ for CIFAR-10 and $77.85\%$ out of $83.64\%$ for MNLI, and with $50\%$ it approaches the original test accuracy of the model, i.e. $92.09\%$ for CIFAR-10 and $81.94\%$ for MNLI. An attacker with access to more than $50\%$ of the data can extract the victim model successfully.
While the risk such an attacker poses is quite high, in practice, if the attacker has access to most of the victim training data, it could have already trained the model on its own, and at most the model serves as a labeling oracle. To better demonstrate this, in Figure 2 we compare the attack accuracy, using soft-label access to the victim model, to the accuracy the attacker would have obtained using real (ground truth) labels. Evidently, the attacker does not benefit much from attacking the model. We further demonstrate this point by comparing the attack accuracy when the attacker has label-only access. We observe that this case is equivalent to the real-label performance, indicating that, besides serving as a labeling oracle for the attacker, the extraction does not show much gain over simply training from scratch. This phenomena aligns with the prior literature (Orekondy et al., 2019b).
4.3 Can surrogate dataset aid extraction?
We now investigate the ME performance gains from the use of additional queries sampled from data distributions that differ from the victim model’s training distribution, yet are ‘similar’ to it, following Truong et al.; Krishna et al.. Here, we consider three adversaries: the first assumes no prior knowledge; the second has partial knowledge about the (related) data; the third has unrestricted access to label any given data point. Namely, for CIFAR-10, we additionally use: Random - uniformly sampled random images; Surrogate - real images sampled from a surrogate (SVHN and CIFAR-100) dataset; DFME - synthetic queries generated using the DFME attack. For Indoor67, CUBS200, and Caltech256, we evaluate only the surrogate variant using the ImageNet dataset, as described by Orekondy et al.. Results are presented in Appendix G. For MNLI we use: Random - nonsensical sequence of random words built from the WikiText-103 corpus (Merity et al., 2017), sampled letter-by-letter; Nonsensical - nonsensical sequence of real words built from the WikiText-103 corpus, sampled word-by-word; Wiki - real sentences sampled from the WikiText-103 corpus.
We measure the improvement in attack accuracy as a function of the additional query budget and prior knowledge. We focus on the attacker with less than $50\%$ IND data access, since in Section 4.2 we show that higher prior knowledge renders attacks unnecessary. Due to the high computational cost of DFME, we only evaluate this type of queries for the attacker with up to $30\%$ prior knowledge.
We first investigate the case where the query budget is constant and is set to the size of the original training dataset, i.e. the maximal query complexity still considered a success. For each prior knowledge percentage $x\%$, we fill the remaining queries with any of the previously described methods. We include the case where all the queries are sampled from the query distribution, and the attacker has no ($0\%$) prior knowledge, to better visualize the effect of the prior knowledge. Note that DFME requires a query budget significantly higher than the size of the training set, so here we set it to 20M.
Refer to caption (a) CIFAR-10 Refer to caption (b) MNLI
@@ 4.3 Can surrogate dataset aid extraction?
Figure 3: Effect of augmenting the attacker’s queries with additional queries sampled from other data distributions, compared to our baseline attacker, which only uses its prior knowledge. We fix the query budget to be the size of the original training set for a fair comparison. Attackers with more prior knowledge over the true distribution do not benefit much by augmenting the query set. Moreover, as previously shown by Truong et al., the similarity between the victim’s data distribution and the query distribution has a large impact on the attack’s success.
The results, presented in Figure 3, show that an attacker with more prior knowledge gains little benefit from additional queries. In the case of CIFAR-10, an attacker with $50\%$ prior knowledge only increases its success rate by up to $1.4\%$. For MNLI, this number is $0.6\%$. We do observe an improvement for the low-query setting, yet it is not enough to successfully extract the victim model – with the exception of the CIFAR-100 and the DFME cases. CIFAR-10 and CIFAR-100 are very similar datasets, with some classes being semantically similar e.g. automobiles and trucks. This is not the case for DFME, where the queries are synthesized, however, in this case, a very large query budget (20M) is used. The high budget can be explained by the intuition provided in Section 3, where in the absence of knowledge of the relevant regions to sample queries from, the adversary is required to traverse the input space and use a large query budget for this. In Appendix C we evaluate the effect of additional queries for a given prior knowledge percentage, showing diminishing returns from them.
We thus conclude that ME attacks need a formal definition of their adversarial goals – we observe that the adversary can achieve its goal of training an accurate model with just IND datapoints labeled, without extra queries to the victim model. We find that in all our experiments with different attacks, the victim model acts as a labeling oracle, and the adversary simply trains with labeled data.
5 Limiting Out-Of-Distribution Effectiveness
@@
Figure 4: The effect of applying our OOD component with different values of $\tau$ against an attacker that utilizes additional queries. In all cases other than DFME, the attacker adds $\|D_{train}\|$ additional queries, and for DFME adds 20M queries. When comparing the results to the original setting (real model), where the OOD region is unmodified, we can see a clear decrease in the attack accuracy.
Following our motivating example in Section 3, we explore the extent to which the attacker relies on an implicit assumption that the IND decision boundaries can be inferred from the OOD ones. If this was not the case, then an intelligent attacker would have to distinguish between IND and OOD samples, or otherwise waste queries and potentially learn non-existent decision boundaries. We augment our classifiers with OOD detection components. Similarly to the Figure 1(b-c), we modify the decision boundaries in the OOD region such that the adversary is unable to statistically tell IND or OOD responses apart. In this setting, we observe that attacks are no longer query effective (unless they have prior knowledge of the data distribution) because not all queries provide an informative response—and it is not easy for the adversary to tell which queries do.
5.1 Sampling complexity intuition
5.2 On hardness of OOD detection
By modeling the attacker’s prior knowledge over the data distribution as a function of the distance between the victim’s data distribution and the attacker’s query distribution, we can see that the probability of sampling an informative IND datapoint decreases when this distance gets larger. As such, the attacker is inevitably sampling from both IND and OOD regions. Following the previously described assumption where the task-related information is limited to IND queries, the attack success is now a function of the ability to solve a non-trivial OOD detection task: an adversary who is unable to distinguish between the IND and OOD samples inevitably faces a higher query complexity for a successful attack. Despite significant progress in OOD detection, the literature still remains largely empirical (Fang et al., ; Hendrycks and Gimpel, 2016b; Carlini and Wagner, 2017) and finds that OOD detection is not at all trivial. Interestingly, robust OOD detection is linked to research into model robustness. Recently, Tramer demonstrated that robust adversarial example detection is as hard as classification. Here, robustness of the detector refers to consistent detection over the $\epsilon$-radius around a data point, representing the maximal distance for an adversarial example. Fundamentally, Tramer demonstrates that robust detection over an $\epsilon$ region around a given (not necessarily adversarial) data point implies, albeit inefficient, an ability to successfully classify for at least $\epsilon/2$ region.
Simply put into Bayesian interpretation, Tramer says that an ability to compute $\mathbb{P}_{\mathcal{D}}(x=x_{i},y=y_{i})$, i.e. perform OOD detection robustly with knowledge of the labels, implies an ability to classify robustly $\mathbb{P}_{\mathcal{D}}(y=y_{i}|x=x_{i})$. Intuitively, that follows from the Bayes rule with $\mathbb{P}_{\mathcal{D}}(x=x_{i},y=y_{i})=\mathbb{P}_{\mathcal{D}}(y=y_{i}|x=x_{i})*\mathbb{P}_{\mathcal{D}}(x=x_{i})$, where $\mathbb{P}_{\mathcal{D}}(x=x_{i})$ represents prior knowledge over the true data distribution $\mathcal{D}$, i.e. the likelihood of $x_{i}$ coming from $\mathcal{D}$. In this paper we do not make assumptions about the labels – one can certainly imagine an attacker who has access to some labeled data. Do note however that label information is not necessary for a classic out-of-distribution detection i.e. checking if a given point comes from a given distribution $\mathbb{P}_{\mathcal{D}}(x=x_{i})$. There are two possibilities. First, if the attacker knows the labels of the points being queried, then the setting is exactly the same as the one considered by Tramer i.e. robust OOD detection is reducible to robust classification. Second, if the attacker is assumed to be capable of estimating $\mathbb{P}_{\mathcal{D}}(x=x_{i})$ robustly, reduction to classification without labels is impossible – the attacker would simply need to use the model as a labeling oracle.
Practically, this implies that with the OOD component in place, ME might not be the most efficient way for the attacker to achieve their objectives, given the capabilities they have access to. If they have the labels and an OOD detector, then they can already perform classification. If they do not have the labels, they can at most get the benefit of labeling their dataset, after which they get classification. If they are unable to solve the OOD detection problem, they would have to first build an OOD detector (which means collecting costly datapoints), since otherwise OOD queries provide them no information about the IND region. In other words, once we take into account the OOD component in place, a model extraction attacker is only as good as their knowledge of the underlying dataset; but if this knowledge includes labels, often there is no need to extract models in the first place.
5.3 Out-of-Distribution instrumentation
In Sections 5.1 and 5.2, we discussed the intuition behind our hypothesis that most previous attacks make an implicit assumption about the usefulness of OOD queries. We showed that when this assumption does not hold, and the OOD region is not indicative of the IND behaviour, ME attack success reduces to mostly being a function of the level of prior knowledge, or an attacker’s ability to solve a non-trivial OOD detection task. We also argued that at a point when an adversary has the ability to solve OOD task, the attacker has enough knowledge to train the model themselves, indicating that the victim model merely serves as a labeling oracle.
In this section we simulate a setting where this implicit assumption does not hold, and empirically show how the attack success is, again, only dependent on prior knowledge of the data distribution. For this evaluation, we revisit the attacker that utilizes both prior knowledge and additional OOD queries, described in Section 4.3. We now propose a simple modification to the victim model that targets the assumption over the OOD region and limits the informativeness of the OOD behaviour.
5.3.1 OOD Implementation Details
We propose a simple method for invalidating the informativeness of the OOD data region. Given the original victim model $\mathcal{V}_{o}$, we create a hybrid victim model $\mathcal{V}_{h}$ by combining $\mathcal{V}_{o}$ with an additional module $\mathcal{V}_{f}$ with different, or additional, decision boundaries. This additional model will be used to provide predictions for OOD queries that differ from the predictions they would have gotten from the original model. For each query $x$, the hybrid model $\mathcal{V}_{h}$ applies some decision rule $R$ to classify $x$ as IND or OOD, and uses the corresponding model for prediction.
We design the additional model $\mathcal{V}_{f}$ such that the decision boundaries of both models would have similar smoothness properties; thus, learning the “fake” boundaries is expected to be of the same level of difficulty, with nearly statistically indistinguishable output distributions. We discuss this further in Section F.4.
We now describe the implementation of the decision rule $R$ and the complex model $\mathcal{V}_{f}$, and expand on details in Appendix F. For the decision rule $R$, we apply some pre-defined threshold $\tau$ over the prediction confidence of $\mathcal{V}_{o}$, with a softmax temperature of $2$ to better calibrate $R$. If the confidence is higher than $\tau$, the hybrid model returns $\mathcal{V}_{o}(x)$, otherwise it returns $\mathcal{V}_{f}(x)$. In Section F.3, we discuss the effect of $\tau$ in more detail. Prediction confidence serves as a naive OOD detector (Hendrycks and Gimpel, 2016a; DeVries and Taylor, 2018) and represents one of the simplest settings for the attacker. More advanced OOD detectors will strictly improve separation IND and OOD queries, therefore decreasing the number of false negative queries, i.e. OOD queries that are classified IND and are provided with a meaningful prediction by $\mathcal{V}_{o}$.
We provide the full description and details of the fake model $\mathcal{V}_{f}$ in Appendix F, and only give here a brief overview. We implement $\mathcal{V}_{f}$ by fitting a Gaussian Mixture Model (GMM) to the logits of each class, using the real training data samples. We additionally create anchor points $\mathcal{A}_{c}^{1},\ldots,\mathcal{A}_{c}^{m}$ for each class, using feature representations of each class’ data samples. These will be used to “assign” queries to one of the GMMs for prediction. The anchor points are permuted between the classes, to ensure that tail-of-distribution queries will not be “correctly” assigned to their real class, and therefore leak IND information. We denote this permutation by $\pi_{j}:C\rightarrow C$, for $j\in[m]$.
For a given query sample $x$, we compute its feature representation and find the nearest $L_{2}$ anchor point $\mathcal{A}_{i}^{j}$. We sample fake logits $\tilde{y}$ from the GMM for the chosen anchor point, according to its permutation $i^{\prime}=\pi_{j}[i]$. We return it as the fake model prediction $\mathcal{V}_{f}(x)=\tilde{y}$. This mechanism results with OOD samples being “wrongly” predicted while still exhibiting similar smoothness properties.
5.3.2 OOD and the additional queries
@@
We show how our OOD component reduces the informativeness of OOD queries and thus the accuracy the attacker gained by utilizing additional OOD queries. We evaluate different threshold values ($\tau$) against an attacker that, in addition to its prior knowledge samples, uses additional $\|D_{train}\|$ queries, i.e. the size of the original training dataset. For DFME, we allow that attacker to use 20M additional queries. Different threshold values influence the results by changing the number of additional queries that are predicted by $\mathcal{V}_{f}$ versus $\mathcal{V}_{o}$ – the lower the threshold, the better for the attacker. We measure the false-positive rate (FPR) for the different threshold values in Section F.3.
@@
We present the results of adding our OOD component in Figure 4. The results clearly demonstrate the described effect. As the attacker has less prior knowledge, it is more reliant on the benefit of the additional queries, and, as such, it is more affected by adding our OOD complex model $\mathcal{V}_{f}$. The attack accuracy is thus reduced closer to, or bellow, the accuracy of the baseline (which makes no additional queries). The effect of the OOD component is weaker for some of the settings in our NLP task, specifically the nonsensical and wiki queries. This is due to the similarity between the true data distribution and these query distribution. We discuss this observation in detail in Appendix H.

### v3

@@ 5 Evaluation
To answer the questions described above, we evaluate a range of ME attacks on common vision and language benchmarks. We measure the attacker’s performance as the accuracy difference between the victim and the attacker on the original task. We define a successful attack as one that produces a model that performs on a task nearly as well as the victim model. Unless stated otherwise, in all our experiments we assume that the attacker has a black-box access to the victim model: the adversary issues a query, and the victim model responds with a vector that indicates the probability of classifying the input in each of the task classes. We note that this setting provides the most information to the adversary. In other words, it advantages the ME adversary.
We define by baseline attacker an attacker that has access to a randomly sampled subset of the victim’s true training dataset and uses this data only to query the victim and train the attack model. Since ME quality is mainly affected by the quality of the attacker queries, real victim training data represents one of the best query sets for the attacker (compressed datasets e.g. Wang et al. (2018) or disjoint IND data can potentially lead to an equally capable extraction set). Throughout our work, we quantify different levels of prior knowledge by the percentage of the data available to the attacker and evaluate the attack accuracy as a function of this measure. We also evaluate ME adversaries that utilize a mixture of IND and OOD data.
@@
(a) CIFAR-10
(b) MNLI
@@
Figure 2: A comparison between the baseline attacker, which only uses its prior knowledge, and an attacker that can augment its queries with additional queries sampled from other data distributions. We fix the query budget to be the size of the original training set for a fair comparison. Attackers with more prior knowledge do not benefit much by augmenting the query set.
5.1 Does model extraction work?
We start off by answering the foundational question – does ME work? Our answer to this is yes, ME attacks definitely work, and it is indeed possible to approximate a blackbox model from just queries, matching performance to the victim model. As seen in Figure 2, and as evident in the large number of works in this field, there are many settings in which a ME adversary can obtain a model that has a desirable performance over the task of interest.
However, not all queries are equally useful for ME. To demonstrate this, we evaluate a ME attacks on real models that use queries sampled from data distributions that differ from the victim model’s training distribution, i.e. OOD queries. Here, we consider a number of possible distributions, where some are closer to the true distribution than others. Namely, for CIFAR-10, we use: Random - uniformly sampled random images; Surrogate - real images sampled from a surrogate (SVHN and CIFAR-100) dataset; DFME - synthetic queries generated using the DFME attack. For Indoor67, CUBS200, and Caltech256, we evaluate only the surrogate variant using the ImageNet dataset, as described by Orekondy et al.. Results are presented in Appendix G. For MNLI we use: Random - nonsensical sequence of random words built from the WikiText-103 corpus (Merity et al., 2017), sampled letter-by-letter; Nonsensical - nonsensical sequence of real words built from the WikiText-103 corpus, sampled word-by-word; Wiki - real sentences sampled from the WikiText-103 corpus.
We investigate the case where the query budget is constant and is set to the size of the original training dataset, i.e. the maximal query complexity still considered to be cost-effective. For each prior knowledge percentage $x\%$, we fill the remaining queries with any of the previously described alternative data sources. We evaluate for up to $50\%$ prior knowledge, because we find that the attacker’s query set is dominated by IND samples for higher proportions and is thus less informative. We include the case where all queries are sampled from the query distribution, and the attacker has no ($0\%$) prior knowledge. Note that DFME requires a significantly higher query budget and incurs a high computational cost, therefore we evaluate this setting for a query budget of 20M and with up to $30\%$ prior knowledge.
As can be seen in Figure 2 in the no prior knowledge setting ($0\%$), most attacks were able to obtain a success rate significantly higher than random guess. It is interesting to observe the behaviour of the synthetic queries. The random images or random text based attacks are significantly outperformed by other methods. This comes in contrast to the success of DFME’s synthetic samples performance. This contrast suggests that it is a question of query complexity, where in cases where the samples differ from natural samples, a high query budget must be used (DFME’s success dramatically reduces when lowering the query budget (Lin et al., 2023)). Additionally, it is interesting to observe the margin between the success of different OOD-based attacks and the success of the baseline attacker. As can be seen, for lower prior knowledge settings, most attacks improve upon the baseline attacker, enabling the attacker to learn additional information using the OOD queries. For example, for CIFAR-10, an attacker with $1\%$ prior knowledge can improve its performance by up to $50\%$. However, with higher levels of prior knowledge, the attacker gains little benefit from the additional OOD queries. A CIFAR-10 attacker with $50\%$ prior knowledge only increases its success rate by up to $1.4\%$. In Appendix B we evaluate the effect of additional queries for a given prior knowledge percentage, showing diminishing returns from them.
5.2 Can ME be used with only a few queries?
(a) CIFAR-10
(b) MNLI
Figure 3: Evaluation of the risk posed by an attacker with some prior knowledge over the true data distribution, using different labeling sources. As can be seen, labels provided by the victim model, either in the richer soft-label setting or in the more restrictive label-only setting, does not provide benefit over the real ground truth labels. This shows that the victim is essentially a labeling oracle.
ME seeks to obtain a copy of the victim model while incurring reduced costs compared to training from scratch. In Section 3.2 we note that one way to reduce costs is to limit the interaction with the victim model to the minimum – that is, reduce the overall ME query budget. In this subsection we investigate such reduction and discover that it is indeed sometimes possible. In Figure 2 we demonstrate that limiting the query budget for all cases but DFME, still leads to the high attack performance. What is more, even the baseline attacker with only $10\%$ of the data, i.e. reducing the query complexity by a factor of 10, manages to develop a model with a relatively high performance, i.e. $79.21\%$ compared to $95.54\%$ for CIFAR-10 and $77.85\%$ compared to $83.64\%$ for MNLI; with $50\%$ it approaches the original test accuracy of the model, i.e. $92.09\%$ for CIFAR-10 and $81.94\%$ for MNLI. We find that the baseline attacker with more than $50\%$ of the original victim training set can extract the model (see Figure 3). Although we observe this phenomenon consistently in the case above, in our answer to the next question we will show that such good performance heavily depends on the prior IND knowledge or requires OOD data that, as described in Section 4, informs the attacker about the IND performance.
5.3 Can ME be used to reduce data costs?
@@
Figure 4: The effect of controlling OOD informativeness with different values of $\tau$ against an attacker that utilizes additional queries. In all cases other than DFME, the attacker adds the size of the training set additional queries, and for DFME adds 20M queries. When comparing the results to the original setting (real model), where the OOD region is unmodified, we can see a clear decrease in the attack accuracy.
A common threat model assumes that IND is scarce and expensive. As such, it is desirable to reduce data collection costs by avoiding, or minimizing the use of IND queries, and prioritizing the use of OOD queries. We find that in some cases ME attacks can indeed do so. More precisely, in Figure 2 we show that OOD queries provide a feasible substitute for IND queries in lower prior knowledge settings. However, we claim that this is only possible because OOD queries inform the adversary about the IND performance i.e. IND decision boundaries can be inferred from the OOD ones, either because the distributions are sufficiently similar or because the model optima are predictable. To explicitly estimate if all OOD queries will aid ME performance, we instrument the model with the OOD detector as described in Section 4.3. Our instrumentation, in essence, allows us to control how much information can be derived about the IND from the OOD region.
Per Section 4, an intelligent attacker would have to distinguish between IND and OOD samples to bypass the instrumentation or otherwise waste queries and potentially learn non-existent decision boundaries. A poor attack policy would inevitably learn fake decision boundaries, while a strong attacker would learn to avoid them. Theoretically, in Section 4.1 we discuss that with no prior knowledge, it should be improbable for an attacker to sample from the IND region alone. Furthermore, in Section 4.2 we argue that when samples from both regions are considered, prior knowledge is required to distinguish between the two.
@@
We empirically evaluate this argument using the methodology described in Section 4.3. We show that when the informativeness of the OOD region is deteriorated, OOD queries cannot replace the need for expensive IND queries. We evaluate different threshold values ($\tau$) against an attacker that, in addition to its prior knowledge samples, uses additional $\|D_{train}\|$ queries, i.e. the size of the original training dataset. For DFME, we allow that attacker to use 20M additional queries. Different threshold values influence the results by changing the number of additional queries that are predicted by $\mathcal{V}_{f}$ versus $\mathcal{V}_{o}$ – the lower the threshold, the better for the attacker. We measure the false-positive rate (FPR) for the different threshold values in Section F.3.
@@
The results in Figure 4 clearly demonstrate the described effect. As the attacker has less prior knowledge, it is more reliant on the benefit of the additional queries. As such, it is more impacted by the fake boundaries that control OOD informativeness. The attack accuracy is thus reduced closer to, or below, the accuracy of the baseline, which makes no additional queries. This effect is weaker for some of the settings in our NLP task, specifically the nonsensical and wiki queries. This is due to the similarity between the true data distribution and these query distributions. We discuss this observation in more detail in Appendix H.
5.4 Can ME be used both to reduce the data costs and use only a few queries?
Both Q2 and Q3 assessed that ME costs could be reduced by either using substitute data or reducing the number of queries; but can an adversary achieve both? The discussion of Q3 gives a clear answer to this question - no.
We showed that unless the OOD dataset is chosen in such a way as to inform the attacker about the IND behavior, the attacker must either rely on the IND data available to them or utilize a very large query budget. It is impossible to achieve both in this setting.
5.5 Can ME be used to reduce data labeling costs?
In Q2, Q3, and Q4, we discussed ME assuming a reduction of costs on data collection and the overall number of queries. We concluded that in generality, such reduction is unattainable, and OOD queries do not trivially replace IND queries. In this section, we turn to label cost reduction.
Label cost reduction could be achieved in two ways: by reducing the cost of a single query or, alternatively, by increasing the amount of information that a single query provides. Note that an increase in the former does not always mean that overall cost increases, as combined with the increase in the latter, they can cost less overall.
We find that in the classification setting considered in this work, we could not find a method to consistently increase the informativeness of the response, suggesting that the victim model merely serves as a labeling oracle, even in cases where seemingly more information is provided than just a label. To demonstrate this, we revisit our baseline attacker and compare its performance when trained on: 1) soft-label i.e. full probability vector; 2) using label-only; and finally, 3) using the real ground truth labels. Results presented in Figure 3 demonstrate that the attacker’s performance is nearly equivalent in all three cases. This phenomenon aligns with the prior literature (Orekondy et al., 2019b). Additionally, we examined each of the cases from a computational perspective and compared their convergence rates. In Figure 5 we present this comparison for the case where the attacker utilizes $30\%$ prior knowledge. We observe that the convergence rate is similar across the three attackers, i.e. the attacker does not even “learn faster” by querying the victim model. In Appendix C we additionally show this for the $5\%$ and $60\%$ prior knowledge settings and observe a similar phenomenon.
We additionally examine the response informativeness effect for attackers who can utilize additional queries for other query distributions. Results, presented in Appendix D, show that for lower prior knowledge settings, and for query distributions that significantly differ from the true distribution, soft-labels provide some benefit over label-only access. However, this effect diminishes as the attacker uses more prior knowledge, or acquires access to a query distribution which is closer to the true distribution.
(a) 30% CIFAR-10
(b) 30% MNLI
In light of the finding that the victim model serves as a labeling oracle, how cost-effective is querying the victim? We discussed current annotations costs in Section 3.2, where we suggested that for many use cases, including current ME benchmarks, crowdsourcing annotations may be a cost effective alternative. However, this is not true for all cases, as e.g. medical data will most likely have higher data collection and labeling costs. We therefore conclude that it is sometimes possible to reduce labeling costs by not running ME and instead finding an alternative labeling provider.